OpenStudy (anonymous):

Joan has $1.45 in dimes and nickels. She has 20 coins altogether. How many coins of each kind does she have?

OpenStudy (anonymous):

Lets let d = the number of dimes and n = the number of nickels. Are you learning systems of equations where you would have 2 equations each with 2 variables and you need to solve for both variables? If so, what method of solving are you using? Ellimination (or addition) or Substitution?

OpenStudy (anonymous):

9 + 5.5

OpenStudy (anonymous):

I use elimination Is this right N+d=20 5n+10d=145

OpenStudy (anonymous):

You can make this simpler by understanding the concept of guess and check. Though I have developed a new method but is tedious... to list down.

OpenStudy (anonymous):

yes! Most instructors will not accept guess and check if they are asking for an algebraic solution.

OpenStudy (anonymous):

actually, the second equation is 5n + 10d = 1.45 Now do you know how to solve the system?

OpenStudy (anonymous):

Yes i know now, i got n=11 and d=9

OpenStudy (anonymous):

if u use 5n+5d is not equals to 1.45 instead you use 11nickel + 9dimes = 1.45